Proposal #1 would extend trade credit to some customers that previously have been denied credit because they were considered poor risks.   Sales are projected to increase by $240,000 per year if credit is extended to these new customers.  Of the new accounts receivable generated, 6% are projected to be uncollectible.  Additional collection costs are projected to be 2% of incremental sales (whether they actually end up collected or not), and production and selling costs are projected to be 78% of sales.  Your firm expects to pay a total of 30% of its income after expenses in taxes.

 

1)     Compute the incremental income after taxes that would result from these projections:

 

2)     Compute the incremental Return on Sales if these new credit customers are accepted:

 

If the receivable turnover ratio is expected to be 3 to 1 and no other asset buildup is needed to serve the new customers…

3)     Compute the  additional investment in Accounts Receivable

4)     Compute the incremental Return on New  Investment

 

5)     If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers?  Explain.
Answer #1

1. Incremental Annual Income after Taxes= $23520


2.Incremental Return on sales=Incremental Annual Income after Taxes/Incremental Sales


=23,520/240,000=9.8%


3. Additional Investment in Account receivable=$80,000


4. Incremental return on New investment=Incremental Annual Income after Taxes/Additional Investment


=23,520/80,000=29.4%


5. Yes, the number suggest that trade credit should be extended to these new customers because the Incremental return on new investment is 29.4%, which is greater than the 20% rate of return on investment.


answered by: Mr. B